A pharmacist wishes to prepare 2 liters of 85% v/v of Ethanol. Concentrations available in the pharmacy include 50% and 95%. How much of each solution, ethanol and distilled water, would be needed to arrive to the desired amount and concentration?

Please help! I am confused with the water bit... is dilution involved? Many thanks in advance.

Answer by Alan3354(69443)

Do it like this one. Only the numbers are different.
-----------
Amanda wants to make 8 gal. of a 20% saline solution by mixing together a 56% saline solution and a 8% saline solution. How much of each solution must she use?
================
e = amount of 8%
f = amount of 56%
---
e+ f = 8 (total solution)
8e + 56f = 20*8 (total saline)
---
e+ f = 8
e + 7f = 20
------------ Subtract
-6f = -12
f = 2
e = 6